# AltGraph: Redesigning Quantum Circuits Using Generative Graph Models for Efficient Optimization

### 23 Feb 2024 v1

@@ Abstract
I Introduction
@@
There are two methods of performing quantum circuit transformation (referred to as circuit transformation for brevity) [1] namely, rule-based and search-based. The former uses gate-cancellation rules to optimize quantum circuits[2], while the latter uses machine learning techniques to find functionally equivalent quantum circuits from the search space of quantum circuits [1, 3, 4, 5, 6]. These techniques primarily use Reinforcement Learning (RL) methods e.g., [7] uses an RL agent to determine the correct set of gate cancellations to reach optimal depth on ICM+H circuits. Since ICM+H circuits contain only CNOT and Hadamard gates which are easy to optimize, this technique may not necessarily work on quantum circuits involving other gates. [4] overcomes this issue by using parametric quantum circuits but require circuits to be hardware compliant prior to performing optimization. [5] uses generative methods by generating circuits as Quantum Assembly Language (QASM) instructions, however the search space of the generation algorithm is limited and also does not sample new circuit representations that may fundamentally address the issues like hardware constraint. We address these shortcomings by proposing AltGraph: usage of generative graph models that produce an equivalent Directed Acyclic Graph (DAG) of the original quantum circuit DAG from which we reconstruct the corresponding quantum circuit. We perturb the latent space of the models to generate equivalent quantum circuits some of which may be more compatible with the hardware coupling map and/or enable better optimization leading to compact circuits. We use multiple graph generation models, specifically DAG Variational Autoencoder (D-VAE) Gated Recurrent Unit (GRU), D-VAE Graph Convolutional Network (GCN) and Deep Generative Model for Graphs (DeepGMG) and perform comparative analysis.
@@
The rest of the paper’s organization is as follows: Section II illustrates the background information required to understand the paper and related work. Section  III presents the approach to convert the models to valid quantum circuit graph generators. Section  IV contains visualizations of each model’s abilities to process quantum circuits. Section  V analyzes the results, and Section  VI summarizes the work.
II Background and Related Works
II-A Fundamentals of Quantum Computing
II-A1 Qubits
@@
II-A2 Quantum Gates
Quantum gates, which are unitary matrix functions, modify the states of one or multiple qubits. Dependent on qubit technology, gates are implemented in various ways. For instance, superconducting qubits utilize microwave pulses [8], while trapped ion and quantum dots employ laser pulses [9, 10]. On the other hand, Nuclear Magnetic Resonance (NMR) qubits use radio frequency pulses. The speed at which these gates operate can differ significantly based on the technology, with photonic qubits working in picoseconds and NMR qubits taking up to several seconds [11]. Gates that affect a single qubit encompass the X (NOT) gate, H (Hadamard) gate, and rotational gates like $R_{X}$/$R_{Y}$/$R_{Z}$, and the $U$ gate. Gates impacting multiple qubits consist of the CNOT (controlled-NOT) gate, Toffoli gate, controlled rotation gates, and the Peres gate [12].
II-A3 Quantum circuit
A quantum circuit is a structured series of gate tasks carried out chronologically. It starts with setting up the initial state of the qubits, known as state initialization. Following this, gate operations within the circuit modify qubits to achieve a specific state. The process concludes with a measurement using a particular gate. These quantum activities occur in a quantum Hilbert space [13].
II-A4 Transpilation
@@
II-B Quantum Circuit Transformation
There are two types of circuit transformation: rule-based approach and search-based approach. A rule-based approach involves utilizing gate-cancellation rules on particular subcircuits to reduce them into smaller amounts of gates with lower depths. For example, two back-to-back Pauli X gates computationally equate to identity operation (I) as $\begin{bmatrix}0&1\\ 1&0\end{bmatrix}\begin{bmatrix}0&1\\ 1&0\end{bmatrix}=\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$. Similar behavior exists for other single-qubit gates and multi-qubit gates. This kind of circuit transformation is typically incorporated in quantum compilers like Qiskit [2]. Search-based approaches offer flexibility in terms of optimization. Rather than relying on gate-cancellation rules, this approach searches for a functionally equivalent circuit. Typically, Machine Learning (ML) techniques such as Reinforcement Learning (RL), generative models, or other methods aid the approach. In RL-based search, the current quantum circuit is the state, and the original version of the circuit is the environment. The circuit transformation performed in each step is the action, and based on some input, the agent performs the most rewarding action. Some RL-based works include \⃝raisebox{-0.9pt}{1} [7], that uses D-VAE to encode a quantum circuit DAG and send that encoding to an RL agent which then decides what action to perform among the possible set of updates, \⃝raisebox{-0.9pt}{2} [1], that uses Hierarchical Advantage Learning (HAL) combined with Proximal Policy Optimization (PPO) [3] to perform reinforcement learning, and \⃝raisebox{-0.9pt}{3} [4] that also uses PPO along with Advantage Actor Critique (AAC) scheme-based agent. In contrast, generative approaches such as [5] generate quantum circuits as a Quantum Assembly Language (QASM) sequence of instructions. Other methods include \⃝raisebox{-0.9pt}{1} [6] uses Monte Carlo tree search framework to generate hardware compliant quantum circuits, and \⃝raisebox{-0.9pt}{2} [14] performs ZX-calculus on graphs to perform circuit transformation.
Figure 1: (a) Sample 4 qubit 16 gate quantum circuit with (b) the DAG rendering of the circuit.
II-C Message Passing
@@ II-C1 Simultaneous
Figure 2: Overview of AltGraph process for generating alternate representation of quantum circuits.
II-C2 D-VAE Asynchronous
@@
II-D Generative Graph Models
Generative graph models create new equivalent models based on input graphs and various constraints. D-VAE encodes a DAG to the latent space, then the encoding is used to obtain reconstructed DAGs. For detailed analysis, different variations of the D-VAE generative graph model [15] are used. Namely, D-VAE GRU, D-VAE GCN, and D-VAE DeepGMG [16]. All three versions use topological ordering to create their DAGs, differing mainly in the encoding method.
II-D1 D-VAE GRU
The standard D-VAE approach uses the Gated Recurrent Unit (GRU), which encodes the current node’s state using the node’s predecessors’ states and the current node type. D-VAE uses asynchronous message passing [17] and topological sorting to ensure all information is available to the node before encoding. D-VAE defaults to bi-directional encoding, where upon completion of forward encoding D-VAE reverses the graph and the node states are updated by repeating the same process.
II-D2 D-VAE GCN
The second D-VAE approach uses GCN. D-VAE allows the designer to decide how many rounds of simultaneous message passing [18] they desire to create the encodings of the nodes. To initialize the encoding D-VAE uses the neighbor (or predecessors if uni-directional) node types to initialize each node’s state. In additional rounds of encoding neighbor (or predecessor) node states are used to update the state. Standard GCN models are primarily for encoding, so D-VAE applies the GRU method for GCN decoding.
II-D3 D-VAE DeepGMG
The third D-VAE approach uses DeepGMG [16]. The original DeepGMG work focuses on undirected graph generation. It also does not use topological ordering and uses a random order instead. To make a more valid comparison, D-VAE updates the design of DeepGMG to use topological ordering, allowing for multiple iterations of message passing and the processing of directed graphs. DeepGMG relies on a similar simultaneous message-passing scheme as GCN with an important update. DeepGMG, instead of only relying on surrounding node information, also accounts for an edge feature vector to better understand node connections.
III AltGraph: Proposed Approach
III-A Approach
@@
III-B Loss
D-VAE used teacher forcing to get faster convergence of models. For loss evaluation, D-VAE used encoder KLD and the model’s ability to predict the correct node. D-VAE also evaluated the loss based on the model’s ability to predict edges using the “true” node. While these features are viable when dealing with graphs and ensuring similarly designed graphs, the goal is to generate equivalent quantum circuits. Structurally similar graphs are not a guarantee of equivalent quantum circuits. It’s necessary to compare density matrices to evaluate quantum circuit equivalence. The loss process is updated to generate the nodes and edges rather than forcing edge predictions based on the “true” node. This forces complete graph reconstruction for each input, allowing density matrix comparison. AltGraph’s loss calculation uses the KLD of the encoder, the mean squared error (MSE) of the original circuit’s density matrix, and the reconstructed circuit’s density matrix.
III-C Graph Generation
@@ III-C1 Number of Qubits
III-C2 Gates & Gate Connections
@@ III-C3 Circuit Types
IV Results
IV-A Setup and Evaluation Framework
In generating equivalent quantum circuits, the maximum gate predictions are capped at the original circuit gate count. This can yield higher loss values but penalizes models for larger circuits, aligning with the aim of more hardware-friendly circuits. AltGraph’s performance is assessed with qubit counts of 2, 4, and 6 and gate counts of 16, 24, and 32 per qubit. A total of 300 random circuits for each size are produced using gates from Section III-C3. The data follows a 90-10 training/testing split. The same 300 samples are used for all models, ensuring consistency. Each model is trained thrice per circuit size, and reported values represent the average outcomes.
TABLE I: Average gate reduction (%) and loss per model per gate count.
@@ IV-A Setup and Evaluation Framework
IV-B Results and Analysis
Table I presents the mean training loss and gate reduction percentages for three models: AltGraph GRU, AltGraph GCN, and AltGraph DeepGMG across various circuit sizes. These percentage reductions are derived from comparing the number of gates between the transpiled original and the transpiled reconstructed circuits.
@@ IV-B Results and Analysis
Figure 3: Averaged gate reduction per circuit size for (a) AltGraph GRU, (b) AltGraph GCN,(c) AltGraph DeepGMG.
Figure 4: Averaged depth reduction per circuit size for (a) AltGraph GRU, (b) AltGraph GCN, (c) AltGraph DeepGMG.
@@
TABLE II: Average gate reduction (%) per model per gate count
@@
TABLE III: Average gate reduction per model
Model Average Gates Reduced Average Reduction%
AltGraph GRU 35.67 37.55
AltGraph GCN 33.78 37.17
AltGraph DeepGMG 32.87 34.73
Comparative Analysis of Model Performances: \⃝raisebox{-0.9pt}{1} From Table II, AltGraph GRU stands out in its gate count reduction capabilities when compared to both GCN and DeepGMG for circuits with 24 and 32 gates. However, for 16-gate circuits, the GCN model surpasses, showing a 5.11% performance difference over the GRU model. The cumulative gate reduction performance of each model is detailed in Table III. \⃝raisebox{-0.9pt}{2} Upon inspecting Table III, it’s evident that the AltGraph GRU, on average, trims more gates than either the GCN or DeepGMG models. Yet, GCN remains a strong contender, lagging behind GRU by a mere 1.02%. The collective depth reduction performance for each model is compiled in Table IV. \⃝raisebox{-0.9pt}{3} A closer look at Table IV reveals that the AltGraph DeepGMG consistently outperforms in depth reduction when set against the GCN and GRU models. Nevertheless, the GRU model isn’t far behind, showing only a slight 5.98% performance gap to the DeepGMG. For a comprehensive view of the average MSE for test circuits, refer to the last column of Table IV. \⃝raisebox{-0.9pt}{4} Finally, analyzing Table IV, it becomes clear that AltGraph GRU excels in achieving a smaller density matrix difference than both the GCN and DeepGMG models.
TABLE IV: Average depth reduction and test circuit MSE per model
Model Average Depth ReductionAverage MSE
AltGraph GRU 37.75 .0074
AltGraph GCN 35.54 .27
AltGraph DeepGMG 40.08 .13
V Discussion
V-A Mixed Optimization Techniques
Drawing from Table II, it’s evident that graph generative models present a promising avenue for gate reductions across different circuit dimensions. However, there’s scope for refining the models further. The circuits used in this study were randomly generated and didn’t undergo optimization using conventional rule-based methodologies. If models were trained on optimized circuits, it’s plausible that there would be a more significant overall reduction between original and reconstructed circuits, as the model would target only substitutions not already addressed. Implementing rule-based optimization post circuit reconstruction could yield added reductions.
V-B MSE Constraints

### 21 Mar 2024 v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June, 2024; San Francisco, CAisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
There are two methods of performing quantum circuit transformation (referred to as circuit transformation for brevity) (Li et al., 2023) namely, rule-based and search-based. The former uses gate-cancellation rules to optimize quantum circuits(Anis et al., 2021), while the latter uses machine learning techniques to find functionally equivalent quantum circuits from the search space of quantum circuits (Li et al., 2023; Schulman et al., 2017; Fösel et al., 2021; Xu et al., 2022; Zhou et al., 2020). These techniques primarily use Reinforcement Learning (RL) methods e.g., (Moflic et al., 2023) uses an RL agent to determine the correct set of gate cancellations to reach optimal depth on ICM+H circuits. Since ICM+H circuits contain only CNOT and Hadamard gates which are easy to optimize, this technique may not necessarily work on quantum circuits involving other gates. (Fösel et al., 2021) overcomes this issue by using parametric quantum circuits but require circuits to be hardware compliant prior to performing optimization. (Xu et al., 2022) uses generative methods by generating circuits as Quantum Assembly Language (QASM) instructions, however the search space of the generation algorithm is limited and also does not sample new circuit representations that may fundamentally address the issues like hardware constraint. We address these shortcomings by proposing AltGraph: usage of generative graph models that produce an equivalent Directed Acyclic Graph (DAG) of the original quantum circuit DAG from which we reconstruct the corresponding quantum circuit. We perturb the latent space of the models to generate equivalent quantum circuits some of which may be more compatible with the hardware coupling map and/or enable better optimization leading to compact circuits. We use multiple graph generation models, specifically DAG Variational Autoencoder (D-VAE) Gated Recurrent Unit (GRU), D-VAE Graph Convolutional Network (GCN) and Deep Generative Model for Graphs (DeepGMG) and perform comparative analysis.
@@
The rest of the paper’s organization is as follows: Section 2 illustrates the background information required to understand the paper and related work. Section  3 presents the approach to convert the models to valid quantum circuit graph generators. Section  4 contains visualizations of each model’s abilities to process quantum circuits. Section  5 analyzes the results, and Section  6 summarizes the work.
2. Background and Related Works
2.1. Fundamentals of Quantum Computing
2.1.1. Qubits
@@
2.1.2. Quantum Gates
Quantum gates, which are unitary matrix functions, modify the states of one or multiple qubits. Dependent on qubit technology, gates are implemented in various ways. For instance, superconducting qubits utilize microwave pulses (Krantz et al., 2019), while trapped ion and quantum dots employ laser pulses (Benhelm et al., 2008; Nguyen et al., 2012). On the other hand, Nuclear Magnetic Resonance (NMR) qubits use radio frequency pulses. The speed at which these gates operate can differ significantly based on the technology, with photonic qubits working in picoseconds and NMR qubits taking up to several seconds (Ladd et al., 2010). Gates that affect a single qubit encompass the X (NOT) gate, H (Hadamard) gate, and rotational gates like $R_{X}$/$R_{Y}$/$R_{Z}$, and the $U$ gate. Gates impacting multiple qubits consist of the CNOT (controlled-NOT) gate, Toffoli gate, controlled rotation gates, and the Peres gate (Peres, 1985).
2.1.3. Quantum circuit
A quantum circuit is a structured series of gate tasks carried out chronologically. It starts with setting up the initial state of the qubits, known as state initialization. Following this, gate operations within the circuit modify qubits to achieve a specific state. The process concludes with a measurement using a particular gate. These quantum activities occur in a quantum Hilbert space (Von Neumann, 2018).
2.1.4. Transpilation
@@
2.2. Quantum Circuit Transformation
There are two types of circuit transformation: rule-based approach and search-based approach. A rule-based approach involves utilizing gate-cancellation rules on particular subcircuits to reduce them into smaller amounts of gates with lower depths. For example, two back-to-back Pauli X gates computationally equate to identity operation (I) as $\begin{bmatrix}0&1\\ 1&0\end{bmatrix}\begin{bmatrix}0&1\\ 1&0\end{bmatrix}=\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$. Similar behavior exists for other single-qubit gates and multi-qubit gates. This kind of circuit transformation is typically incorporated in quantum compilers like Qiskit (Anis et al., 2021). Search-based approaches offer flexibility in terms of optimization. Rather than relying on gate-cancellation rules, this approach searches for a functionally equivalent circuit. Typically, Machine Learning (ML) techniques such as Reinforcement Learning (RL), generative models, or other methods aid the approach. In RL-based search, the current quantum circuit is the state, and the original version of the circuit is the environment. The circuit transformation performed in each step is the action, and based on some input, the agent performs the most rewarding action. Some RL-based works include \⃝raisebox{-0.9pt}{1} (Moflic et al., 2023), that uses D-VAE to encode a quantum circuit DAG and send that encoding to an RL agent which then decides what action to perform among the possible set of updates, \⃝raisebox{-0.9pt}{2} (Li et al., 2023), that uses Hierarchical Advantage Learning (HAL) combined with Proximal Policy Optimization (PPO) (Schulman et al., 2017) to perform reinforcement learning, and \⃝raisebox{-0.9pt}{3} (Fösel et al., 2021) that also uses PPO along with Advantage Actor Critique (AAC) scheme-based agent. In contrast, generative approaches such as (Xu et al., 2022) generate quantum circuits as a Quantum Assembly Language (QASM) sequence of instructions. Other methods include \⃝raisebox{-0.9pt}{1} (Zhou et al., 2020) uses Monte Carlo tree search framework to generate hardware compliant quantum circuits, and \⃝raisebox{-0.9pt}{2} (Duncan et al., 2020) performs ZX-calculus on graphs to perform circuit transformation.
Figure 1. (a) Sample 4 qubit 16 gate quantum circuit with (b) the DAG rendering of the circuit.
2.3. Message Passing
@@ 2.3.1. Simultaneous
Figure 2. Overview of AltGraph process for generating alternate representation of quantum circuits.
2.3.2. D-VAE Asynchronous
@@
2.4. Generative Graph Models
Generative graph models create new equivalent models based on input graphs and various constraints. D-VAE encodes a DAG to the latent space, then the encoding is used to obtain reconstructed DAGs. For detailed analysis, different variations of the D-VAE generative graph model (Zhang et al., 2019) are used. Namely, D-VAE GRU, D-VAE GCN, and D-VAE DeepGMG (Li et al., 2018). All three versions use topological ordering to create their DAGs, differing mainly in the encoding method.
2.4.1. D-VAE GRU
The standard D-VAE approach uses the Gated Recurrent Unit (GRU), which encodes the current node’s state using the node’s predecessors’ states and the current node type. D-VAE uses asynchronous message passing (Li et al., 2015) and topological sorting to ensure all information is available to the node before encoding. D-VAE defaults to bi-directional encoding, where upon completion of forward encoding D-VAE reverses the graph and the node states are updated by repeating the same process.
2.4.2. D-VAE GCN
The second D-VAE approach uses GCN. D-VAE allows the designer to decide how many rounds of simultaneous message passing (Kipf et al., 2016) they desire to create the encodings of the nodes. To initialize the encoding D-VAE uses the neighbor (or predecessors if uni-directional) node types to initialize each node’s state. In additional rounds of encoding neighbor (or predecessor) node states are used to update the state. Standard GCN models are primarily for encoding, so D-VAE applies the GRU method for GCN decoding.
2.4.3. D-VAE DeepGMG
The third D-VAE approach uses DeepGMG (Li et al., 2018). The original DeepGMG work focuses on undirected graph generation. It also does not use topological ordering and uses a random order instead. To make a more valid comparison, D-VAE updates the design of DeepGMG to use topological ordering, allowing for multiple iterations of message passing and the processing of directed graphs. DeepGMG relies on a similar simultaneous message-passing scheme as GCN with an important update. DeepGMG, instead of only relying on surrounding node information, also accounts for an edge feature vector to better understand node connections.
3. AltGraph: Proposed Approach
3.1. Approach
@@
3.2. Loss
D-VAE used teacher forcing to get faster convergence of models. For loss evaluation, D-VAE used encoder KLD and the model’s ability to predict the correct node. D-VAE also evaluated the loss based on the model’s ability to predict edges using the “true” node. While these features are viable when dealing with graphs and ensuring similarly designed graphs, the goal is to generate equivalent quantum circuits. Structurally similar graphs are not a guarantee of equivalent quantum circuits. It’s necessary to compare density matrices to evaluate quantum circuit equivalence. The loss process is updated to generate the nodes and edges rather than forcing edge predictions based on the “true” node. This forces complete graph reconstruction for each input, allowing density matrix comparison. The density matrices are taken from the target hardware post-transpilation using Qiskit’s default optimization level 1. This ensures the circuits are compared based on their real values rather than the ideal state. AltGraph’s loss calculation uses the KLD of the encoder, the mean squared error (MSE) of the original circuit’s density matrix, and the reconstructed circuit’s density matrix.
3.3. Graph Generation
@@ 3.3.1. Number of Qubits
3.3.2. Gates & Gate Connections
@@ 3.3.3. Circuit Types
4. Results
4.1. Setup and Evaluation Framework
In generating equivalent quantum circuits, the maximum gate predictions are capped at the original circuit gate count. This can yield higher loss values but penalizes models for larger circuits, aligning with the aim of more hardware-friendly circuits. AltGraph’s performance is assessed with qubit counts of 2, 4, and 6 and gate counts of 16, 24, and 32 per qubit. A total of 300 random circuits for each size are produced using gates from Section 3.3.3. The data follows a 90-10 training/testing split. The same 300 samples are used for all models, ensuring consistency. Each model is trained thrice per circuit size, and reported values represent the average outcomes.
Table 1. Average gate reduction (%) and loss per model per gate count.
@@ 4.1. Setup and Evaluation Framework
4.2. Results and Analysis
Table 1 presents the mean training loss and gate reduction percentages for three models: AltGraph GRU, AltGraph GCN, and AltGraph DeepGMG across various circuit sizes. These percentage reductions are derived from comparing the number of gates between the transpiled original and the transpiled reconstructed circuits.
@@ 4.2. Results and Analysis
Figure 3. Averaged gate reduction per circuit size for (a) AltGraph GRU, (b) AltGraph GCN,(c) AltGraph DeepGMG.
Figure 4. Averaged depth reduction per circuit size for (a) AltGraph GRU, (b) AltGraph GCN, (c) AltGraph DeepGMG.
@@
Table 2. Average gate reduction (%) per model per gate count
@@
Table 3. Average gate reduction per model
Model Average Gates Reduced Average ReductionAltGraph GRU 35.67 37.55 AltGraph GCN 33.78 37.17 AltGraph DeepGMG 32.87 34.73 Qiskit 26.15 18.65 $\ket{ket}$ 73.68 33.5 Quartz (Xu et al., 2022) 30.1 Quarl (Li et al., 2023) 36.6 Monte Carlo Tree (Zhou et al., 2020) 30.0
Comparative Analysis of Model Performances: \⃝raisebox{-0.9pt}{1} From Table 2, AltGraph GRU stands out in its gate count reduction capabilities when compared to both GCN and DeepGMG for circuits with 24 and 32 gates. However, for 16-gate circuits, the GCN model surpasses, showing a 5.11% performance difference over the GRU model. The cumulative gate reduction performance of each model is detailed in Table 3. \⃝raisebox{-0.9pt}{2} Upon inspecting Table 3, it’s evident that the AltGraph GRU, on average, trims more gates than either the GCN or DeepGMG models. Yet, GCN remains a strong contender, lagging behind GRU by a mere 1.02%. To understand how AltGraph performs compared to state-of-the-art systems we include the results from optimizing the same circuits using two rules based methods, $\ket{ket}$ and Qiskit. We also include the reported average improvements of 3 state-of-the-art systems ran on IBM Qiskit non-parametric circuits. (Xu et al., 2022; Schulman et al., 2017; Zhou et al., 2020).
The collective depth reduction performance for each model is compiled in Table 4. \⃝raisebox{-0.9pt}{3} A closer look at Table 4 reveals that the AltGraph DeepGMG consistently outperforms in depth reduction when set against the GCN and GRU models. Nevertheless, the GRU model isn’t far behind, showing only a slight 5.98% performance gap to the DeepGMG. For a comprehensive view of the average MSE for test circuits, refer to the last column of Table 4. \⃝raisebox{-0.9pt}{4} Finally, analyzing Table 4, it becomes clear that AltGraph GRU excels in achieving a smaller density matrix difference than both the GCN and DeepGMG models.
Table 4. Average depth reduction and test circuit MSE per model
Model Average Depth ReductionAverage MSE AltGraph GRU 37.75 .0074 AltGraph GCN 35.54 .27 AltGraph DeepGMG 40.08 .13
5. Discussion
5.1. Mixed Optimization Techniques
Drawing from Table 2, it’s evident that graph generative models present a promising avenue for gate reductions across different circuit dimensions. However, there’s scope for refining the models further. The circuits used in this study were randomly generated and didn’t undergo optimization using conventional rule-based methodologies. If models were trained on optimized circuits, it’s plausible that there would be a more significant overall reduction between original and reconstructed circuits, as the model would target only substitutions not already addressed. Implementing rule-based optimization post circuit reconstruction could yield added reductions.
5.2. MSE Constraints